OpenStudy (anonymous):

Directions: Find the slope of the line. -2x-8y=24

OpenStudy (anonymous):

slope is -2/8 or .25

OpenStudy (anonymous):

Can you show me how you got that answer?

OpenStudy (anonymous):

-2x-8y=24 then -8y=24+2x

OpenStudy (anonymous):

then divide by -8 to isolate y

OpenStudy (anonymous):

u will get y=3-2/8x

OpenStudy (anonymous):

-2/8 =slope